Solar Co. is currently evaluating a proposal to build a new plant that will manufacture solar panels. The company expects the solar panel line will generate annual sales of $115 million/year (from t = 1 to t = 8). Manufacturing expenses (excluding depreciation) are $49 million/year (from t = 1 to t = 8). Overhead expenses at Solar Co's Colorado head office are unaffected by the project and are expected to remain at $2 million/year. Launching the new high-efficiency solar panels is expected to result in a loss of revenue of $11 million/year for its conventional solar panel division (from t = 1 to t = 8). All numbers are before tax. Net working capital is expected to increase by $12 million today (at t = 0) and will remain at the same level until the project ends. Solar Co. has a corporate tax rate of 35%. What is the total cash flow of this project in year 1 (at t = 1; excluding any CCA tax shields)?
Answer #1

Total cash flow of this project in year 1 =

(Annual sales increased due to launch of solar panels - Manufacturing expense )*(1-tax rate)

Where, Annual sales figure increased due to launch of solar panels =Sales from solar panels - Reduction in sales of conventional solar panels=115-11=$104 million

Note:- overhead expense is not to be considered because it is sunk cost.It will not be affected whether project is taken or not.

Hence,Total cash flow=(104-49)*0.65=$35.75 million